Tim's age is now 225% times what it was ten years ago. How old is Tim now?
Answer by ewatrrr(24785)


Hi
Let x represent Tim's age NOW
Question states
2.25*(x-10) = x
2.25x - 22.5) = x
1.25x = 22.5
x = 22.5/1.25
x = 18, Tim's age NOW